OpenStudy (anonymous):

Write 3^-3=1/27 as a logarithmic form.

OpenStudy (yttrium):

using the property that \[\log _{b}x = y \] is same with \[b^y = x\] hence \[\log _{3}(\frac{ 1 }{ 27 } )= -3\]